# RELIEF: Reinforcement Learning Empowered Graph Feature Prompt Tuning

### v2

@@ A.1. Extend to Node-Level Tasks
Fortunately, prior research has successfully unified various levels of tasks as subgraph-level tasks (Liu et al., 2023a; Sun et al., 2023a). This unification is achieved by generating induced graphs for nodes or edges, which enables feature prompting methods to operate on $k$-hop subgraphs. These subgraphs act as proxies for the target nodes or edges during prompt tuning and downstream tasks evaluation. Here, we examine the capability of RELIEF on node classification tasks, alongside other existing feature prompting approaches.
A.2. Details of Datasets
We adopt Cora, Citeseer and Pubmed (Kipf and Welling, 2017) and Amazon-Co-Buy (Computer and Photo) (Shchur et al., 2018) for node classification. Cora, Citeseer, and Pubmed are citation networks, where each node represents a publication with a sparse bag-of-words feature vector, and edges signify citation links. Computer and Photo are segments of the Amazon co-purchase graph, with nodes representing products, edges indicating frequent co-purchases, node features encoded as bag-of-words from product reviews, and class labels derived from product categories. The statistics of these dataset are summarized in Table 4.
A.3. Details of Pre-training Strategies
Dataset # Nodes # Edges # Features # Classes
Cora 2708 5429 1433 7
CiteSeer 3327 9104 3703 6
PubMed 19717 88648 500 3
Computers 13752 491722 767 10
Photo 7650 238162 745 8
@@ A.3. Details of Pre-training Strategies
A.4. Experimental Settings
Data splitting and preprocessing.
We construct 10-shot node classification experiments by randomly sampling 10 nodes per class, with additional 500 nodes each for validation and testing. Considering the average degree of different datasets, we set varying $k$ for generating k-hop subgraphs, specifically 4, 3, 2, 2, 2 for Cora, Citeseer, Pubmed, Computers and Photo, respectively, to ensure an appropriate number of nodes for induced subgraphs.
Implementation.
We employ a 2-layer GIN (Xu et al., 2018) as the fundamental GNN model, which is widely utilized in previous works (Kipf and Welling, 2017; Shchur et al., 2018; Sun et al., 2022), with a hidden dimension of 128. While more advanced subgraph GNN techniques might yield better performance, as noted in (Fang et al., 2024), we opt to use the basic GIN for processing prompted subgraphs to ensure a fair comparison with GPPT and GraphPrompt, as these two methods do not incorporate subgraph-related designs. To achieve more compact node representations, we use Singular Value Decomposition (SVD) to reduce the initial features of each dataset to 100 dimensions. Cross-entropy loss is used as the reward function. The policy network and the projection head are independently updated using two Adam optimizers. Other hyper-parameters are listed in Table 6. The RL hyper-parameter settings are partly referenced from (Jiang et al., 2021), aligning with its established terminology.
@@ A.4. Experimental Settings
A.5. Performance Results
To evaluate the prompt tuning methods, we conducted five rounds of experiments with different random seeds, assessing performance using accuracy and macro F1-score. As illustrated in Table 5, RELIEF consistently achieves top-tier performance across majority of tasks, frequently ranking either first or as the runner-up, underscoring its applicability and effectiveness. On average, RELIEF outperforms fine-tuning by 2.05% and 1.43% on accuracy and macro f1-score, respectively. Additionally, RELIEF surpasses the second-best method, GPF-plus, by 0.89% in accuracy.
In contrast, GPPT and GraphPrompt exhibit inferior performance in most tasks. A plausible explanation is that their prediction results rely on class prototypes, which are challenging to accurately construct in few-shot scenarios with limited labeled data.
Figure 5. The testing accuracy of RELIEF and GPF-plus with and without feature prompts.
[…]
Figure 5 Fully described in the text.
A.6. Case Study
One notable observation, as demonstrated in Table 5, is that RELIEF’s performance lags behind its usual advantage in other tasks on the Computers dataset with GIN pre-trained by MaskedEdge. We thus conduct a case study to investigate the underlying causes.
Despite employing various widely-used training techniques – such as dynamically reducing the learning rate, altering batch sizes, applying weight decay in the Adam optimizer, and adjusting the dropout ratio – the self-supervised pre-training loss of the GIN model in this case still fails to decrease steadily. This indicates inadequate pre-training of the GNN model.
To investigate the discrepancy between GPF-plus (which performs best under this scenario) and RELIEF, we examine testing accuracy curves with and without incorporating prompts. As depicted in Figure 5, for GPF-plus, the testing accuracy without prompts rises at early epochs, reaching a maximum of 20.0, but then almost monotonically decreases to around 7.0, worse than random guessing in the ten-class classification task on the Computers dataset. Conversely, for RELIEF, although there is a downward trend in the first 50 epochs, with a minimum accuracy of 25.8, the accuracy then rapidly increases to about 75.1 and gradually converges.
We attribute the discrepancy in testing accuracy to different extents of disturbance caused by feature prompts, supported by two aspects. Firstly, we compare the APM (Average Prompt Magnitude) of GPF-plus and RELIEF, which are 0.194 and 0.053, respectively, against the original feature scale of 0.502, computed by the mean of L1-norm of each initial node feature averaged on dimension. Additionally, GPF-plus prompts on all nodes while RELIEF has a PCR (Prompt Coverage Ratio) of 0.61. Secondly, we apply linear probing (Kumar et al., 2022) to this case by solely adjusting the projection head, which is widely used in NLP domain to evaluate the expressiveness of the pre-trained model. The final testing accuracy of linear probing is 68.6, as indicated by a horizontal dotted gray line in Figure 5. When discarding prompts, GPF-plus’s accuracy plummets from 68.6 to 7.0, implying it overrides the initial features with overwhelming prompts, resulting in the pre-trained representations (obtained by directly feeding the original node features into the pre-trained GNN) becoming ineffective. In contrast, RELIEF maintains pre-trained knowledge and further generalizes it, leading to an accuracy increase from 68.6 to 77.6.
This raises the question: Why does GPF-plus excel while RELIEF struggles to show its effectiveness? With an inadequately pre-trained GNN model, which neither provides differentiated representations based on initial features nor adjusts itself during prompt tuning, GPF-plus relies on larger feature prompts to create differences in the input space. As a result, it generates discriminative representations better aligned with training labels via the collaboration between prompts and the projection head.
However, this scenario poses challenges for our RL-based RELIEF. As discussed in Section 3.2, the essence of RL is to map states to actions or to state-value functions. Actors are guided by state-values to explore and update, with state-values computed based on rewards. Thus, capturing the relationships between states and rewards is crucial for a powerful agent. In our scenario, states are constructed on node representations, and rewards are derived from the loss function. However, with the GNN model’s insufficient capacity, it may provide similar representations for samples with different labels and distinct representations for samples of the same class. Consequently, a trajectory (composed of transitions sampled on a single graph during exploration) effective on one graph may fail on another graph with similar representations but different labels. This leads to contradictory transitions, distracting the critic from accurately estimating state-value functions, thereby increasing the difficulty of obtaining the optimal policy.
In summary, GPF-plus tends to overshadow the original input to build a new mapping when faced with an inadequately pre-trained GNN model, while RELIEF performs well by interacting in an environment with stable reward dynamics provided by adequately pre-trained models. Therefore, RELIEF and GPF-plus can excel in complementary settings, where well-pre-trained models are more likely to be encountered and are preferable.
Appendix B More Information on Graph Classification
For datasets and pre-training strategies used in graph classification, we adhere to the training procedures established by pioneering researchers (Hu et al., 2019), with pre-trained GIN models available on their GitHub repository  [^1] https://github.com/snap-stanford/pretrain-gnns/tree/master/chem/model_gin.
B.1. Details of Datasets
Dataset # Molecules # Tasks   Dataset # Molecules # Tasks
BBBP 2039 1   ClinTox 1478 2
Tox21 7831 12   MUV 93087 17
ToxCast 8575 617   HIV 41127 1
SIDER 1427 27   BACE 1513 1
B.2. Details of Pre-training Strategies

### v3

@@ 5.3. Few-shot Node Classification
Given that feature prompting methods are originally designed for graph-level tasks (Fang et al., 2024), adapting them to node-level tasks requires careful consideration. Fortunately, prior research has successfully unified various levels of tasks as subgraph-level tasks (Liu et al., 2023a; Sun et al., 2023a). This unification is achieved by generating induced graphs for nodes or edges, which enables feature prompting methods to operate on $k$-hop subgraphs. These subgraphs act as proxies for the target nodes or edges during prompt tuning and downstream tasks evaluation.
Datasets, Baselines and Pre-training Strategies.
We adopt Cora, Citeseer, Pubmed (Kipf and Welling, 2017) and Amazon-Co-Buy (Computer and Photo) (Shchur et al., 2018) for node classification. To complement the node and graph-level pre-training strategies used in graph classification, here we focus on edge-level pre-training. Therefore, we include all previously mentioned feature prompting methods and additionally involve GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) as baselines, which are tailored to GNNs pre-trained at edge-level and capable of transferring to node classification. For pre-training strategies, GPPT uses masked edge prediction (MaskedEdge), while GraphPrompt employs contrastive learning based on edge connectivity (ContraEdge). Descriptions of the two strategies are presented in Appendix B.
Training Scheme.
We construct 10-shot node classification by randomly sampling 10 nodes per class, with additional 500 nodes each for validation and testing. Considering the average degree of different datasets, we set varying $k$ for generating k-hop subgraphs for each dataset. We employ a 2-layer GIN (Xu et al., 2018) as the fundamental GNN model. Cross-entropy loss is used as the reward function. Five rounds of experiments with different random seeds are conducted, and the performance is assessed via accuracy and macro F1-score. More experimental settings are available in Appendix B.3.
Performance Results.
As illustrated in Table 4, RELIEF consistently achieves top-tier performance across majority of tasks, frequently ranking either first or as the runner-up, underscoring its applicability and effectiveness. On average, RELIEF outperforms fine-tuning by 2.05% and 1.43% on accuracy and macro f1-score, respectively. Additionally, RELIEF surpasses the second-best method, GPF-plus, by 0.89% in accuracy. In contrast, GPPT and GraphPrompt exhibit inferior performance in most tasks. A plausible explanation is that their prediction results rely on class prototypes, which are challenging to accurately construct with limited labeled data.
@@
5.4. Ablation Study
Considering the core of RELIEF lies in its two actors for incorporating prompts, we deactivated either one to create two variants. Specifically, “random_d” randomly samples a valid node to prompt at each step, while “random_c” generates a real-valued vector sampled from a random Gaussian distribution. Additionally, we include “linear probing”, which discards the prompting components and adjusts only the projection head.
Figure 5. Effectiveness of discrete and continuous policies.
[…]
Figure 5 Fully described in the text.
Figure 5 shows the comparisons of RELIEF with its variants on three graph-level and two node-level tasks. We find that deactivating either actor results in performance drops due to the disruption of their coordination. Particularly, the discrete actor plays a crucial role in RELIEF. Intuitively, a random prompt context assigned to a target node can be rectified by prompting on the node again in later steps, given the premise of a powerful discrete actor for locating this node. However, with a random node selection policy, a suboptimal feature prompt is difficult to mitigate by prompting on arbitrary nodes, and is further exacerbated by the message-passing mechanism, which spreads chaotic prompts throughout the graph. In addition, ”linear probing”, which only trains the projection head, exhibits significantly inferior performance, clearly demonstrating the contribution of the inserted feature prompts.
Figure 6. Parameter sensitivity analysis.
[…]
Figure 6 fully described in the text.
5.5. Parameter Analysis
We analyze the critical hyper-parameters in RELIEF, including the maximum feature prompt scale per step $z_{\text{max}}$, the number of sub-policies $l$, and the two penalty factors $\alpha_{d}$ and $\alpha_{c}$ related to policy generalization. Figure 6 illustrates the ROC-AUC on ClinTox and accuracy on Computers, with each hyper-parameter varied over its legal range of values. The observations are as follows:
RELIEF performs consistently well across a broad range of hyper-parameter values, demonstrating its robustness. Specifically, when $z_{\text{max}}$ is too large, the excessive magnitude of prompts introduces unnecessary disturbance, leading to performance degradation. Conversely, setting $z_{\text{max}}$ too small weakens the impact of feature prompting, making it more similar to linear probing. By setting $l$ to 1, which discards the policy generalization technique, results in significant performance decline. While increasing $l$ may improve performance, it also requires maintaining more pairs of actors, leading to computational overhead. For $\alpha_{d}$ and $\alpha_{c}$, excessively large values may result in overly punishment, making the policies too conservative and hindering performance improvements.
Due to space constraints, we relocate the scalability analysis of our RELIEF to Appendix C.
5.6. Why RELIEF works?
Lightweight. RELIEF tends to preserve pre-trained graph knowledge rather than overwhelm it. Larger-scale prompts cause the GNN model to generate unfamiliar representations, disrupting coordination between representations and the projection head, potentially leading to negative rewards. These punishments are applied to the actor, urging it to provide prompts that avoid excessive disturbance to the initial input through policy updates.
6. Conclusion
In this paper, inspired by the marginal effect of increasing prompt token length on performance improvement in LLMs, we propose RELIEF, an RL empowered graph feature prompt tuning method. RELIEF enhances the downstream performance of pre-trained GNN models by incorporating only necessary and lightweight feature prompts to original graphs. To tackle this, we employ an RL algorithm designed for hybrid action spaces and integrated policy generalization techniques. Extensive experiments on graph and node-level tasks demonstrate RELIEF’s effectiveness and data efficiency compared to fine-tuning and other prompt-based methods.
Acknowledgements.
Appendix A Details of Graph-level Tasks
For datasets and pre-training strategies used in graph classification, we adhere to the training procedures in (Hu et al., 2019), with pre-trained GIN models available on their GitHub repository  [^1] https://github.com/snap-stanford/pretrain-gnns/tree/master/chem/model_gin.
A.1. Datasets
Dataset # Molecules # Tasks   Dataset # Molecules # Tasks
BBBP 2039 1   ClinTox 1478 2
Tox21 7831 12   MUV 93087 17
ToxCast 8575 617   HIV 41127 1
SIDER 1427 27   BACE 1513 1
A.2. Pre-training Strategies
A.3. Experimental Settings
To create 50-shot scenarios, we randomly sampled 50 graphs from the training set. This procedure introduced randomness, differing from previous works (Fang et al., 2024; Lee et al., 2024) that selected training data in descending order of scaffold occurrence. Given the limited budget of only 50 samples, prior methods resulted in training data with very few scaffold types, exacerbating overfitting. Consequently, the average ROC-AUC of fine-tuning across tasks under the 50-shot setting reported in previous works (Fang et al., 2024; Lee et al., 2024) was 55.40$\pm$2.13%, slightly above a random guess for binary classification tasks. In contrast, our randomness brings about a dataset with more molecular diversity and also prevents scaffold overlap between training, validation and test sets. As a result, the average ROC-AUC under our data splitting method is 60.92% with a smaller standard deviation of 1.46%.
We employ a 5-layer GIN with hidden dimensions of 300. The policy network and the projection head are independently updated using two Adam optimizers. We assess downstream task performance using the validation early stopping protocol, reporting the test ROC-AUC at the best validation epoch. Additionally, considering that 50-shot may be insufficient for some datasets to ensure stable performance at the onset of training, we determine the best validation epoch after 20 epochs. The complete hyper-parameters for graph classification are detailed listed in Table 7.
Appendix B Details of Node-level Tasks
B.1. Datasets
We adopt Cora, Citeseer, Pubmed (Kipf and Welling, 2017) and Amazon-Co-Buy (Computer and Photo) (Shchur et al., 2018) for node classification. Cora, Citeseer, and Pubmed are citation networks, where each node represents a publication with a sparse bag-of-words feature vector, and edges signify citation links. Computer and Photo are segments of the Amazon co-purchase graph, with nodes representing products, edges indicating frequent co-purchases, node features encoded as bag-of-words from product reviews, and class labels derived from product categories. The statistics of these dataset are summarized in Table 6.
Dataset # Nodes # Edges # Features # Classes
Cora 2708 5429 1433 7
CiteSeer 3327 9104 3703 6
PubMed 19717 88648 500 3
Computers 13752 491722 767 10
Photo 7650 238162 745 8
B.2. Pre-training Strategies